OpenStudy (anonymous):

How do you solve x2 + 8x + 25 = 0

OpenStudy (anonymous):

use the quadratric equation

OpenStudy (anonymous):

you complete the square: (x+a)²=x²+2ax+a²

OpenStudy (anonymous):

so compare with what you have x²+2ax+a² x²+2(4)x+25

OpenStudy (anonymous):

woould it be -4 +/- i

OpenStudy (anonymous):

so a= 4 but we still need the a² so we add and substract x²+2(4)x+ 4² -4²+25=0 = (x+4)² = -9

OpenStudy (anonymous):

x= -4+-3i i get

OpenStudy (anonymous):

to test the solutions you can use http://www.wolframalpha.com/input/?i=x%C2%B2+%2B+8x+%2B+25+%3D+0 just don use it to do the whole whomework! =)

OpenStudy (anonymous):

oh i think you want to use the complete the square method here?

OpenStudy (anonymous):

yep, i think its better than learning the quadratic formula...

OpenStudy (anonymous):

go to khan academy if ou dont know completing the square method

OpenStudy (anonymous):

thank you